Agenda/Minutes of the VIS Meeting on 2016/7/29
2016/7/29 13:30 - @ Room 210
Participants: Takahashi, Ishizaki, Tanaka, Shoda, Mark, Fabian, Aso, Ohishi, Okutomi, Sato, Hirata, Akutsu
- Type-A installation procedure review on the 10th in August.
- Cable assembly: update the schedule chart with more details (Takahashi)
Simulation: implement asymmetry in the model to compute length->yaw couplings (Okutomi)
- Geophone noise is 10 times higher than the theoretical one
- Simulation with this new noise level is necessary
- Discuss with Kawamura-san about Liu-kun's availability (Aso)
- Send a clean suit to Kashiwa (size LL 29)
Type B Past Week Summary
- Mark and Fabian went to Kashiwa to clean up glue residue on the dummy BS and do a fit/function test of Tatsumi-san's modified tooling.
- Most of the remaining glue came off with soaking in acetone, but there is still some in the machining grooves in the Al surface. Probably more work is required.
- A set of cleanroom shoes borrowed from the ATC disintegrated when worn (the plastic had perished from age), which put small pieces of foam all over the floor in Hirose-san's cleanroom and required a lot of clean-up with a sticky roller! Cleanroom suits and shoes for use by Type B team at Mitaka/Kashiwa have been ordered, but will take another week to arrive.
- The flag holding parts are now much easier to place and remove, much reducing the risk of breaking flags after gluing.
- The sapphire and metal prisms from both the first BS gluing and the PR3 hanging test mirror were brought back to Mitaka for ultrasonic cleaning to remove glue, so they can be reused.
- We arranged with Hirose-san to collect the dummy BS and mirror box next Thursday 8/4. (This is conditional on the new clean suits arriving.) Further work will be done at Mitaka.
- Hirata-san arranged to borrow the NAOJ Hi-Ace van for the trip.
- Hirata-san found the SR versions of the screws for immobilizing the IM and IRM. These fit much better than the BS ones (despite being the same design) and will be used for the BS.
- Fabian identified 210 g of the 600 g deficiency in the IM mass as due to SS parts on the outside. However identifying the rest would require time-consuming disassembly of the IM, and it's not worth it because whatever the cause, the fix is to design and add more trim mass.
- Hirata-san measured another filter, and worked on buying various small items.
Mark moved most of the contents of a bunch of older Dropbox and OneDrive shared folders into a new /Subsystems/VIS/TypeB folder on the KAGRA Dropbox.
- Mark worked on a list of considerations for locating the BS rack, in preparation for discussions with Miyakawa-san et al.
- Type B ongoing issues
Akutsu-san is working to get wide-mouth OSEMs (2 assembled; first 4 expected to be assembled and washed by 8/8)
Longer cables for BF->PI and SF->PI.
- Connectors RECEIVED, cables being made.
- SR machining test NEARLY DONE. The part looks good and will be used but the ones for SR3 and SRM will be made from scratch.
- Lower breadboard stuff
- Installation to be incorporated into the procedure document. Mark/Takahashi to discuss.
- Breadboard springs - RECEIVED.
- To-do list during Mark's vacation
- Weigh IM/RM/IRM (Fabian, Hirata) - DONE but anomalies found.
- Pack/ship IM/RM/IRM to Kamioka (Hirata/Fabian) - from 8/1.
- Analyze geophone data (Fabian/Shoda) - data taken by Shoda-san; Fabian to find calibration factors.
- Remove glue from dummy BS face (Fabian) - MOSTLY DONE.
- Reglue flags and prisms (Tatsumi, Fabian) - from Friday 8/5.
- Clean wire for BS and RM (Hirata) - DONE
- Get new winch parts for new wire diameters - design received from Ikenoue-san
- Do fit/function test of winches (Fabian/Hirata)
- Ship winches to Kamioka (Hirata/Fabian)
- Work on BS simulation (Fabian)
- Do Inventor assembly for SR2 in chamber and on assembly frame (Hirata) - ONGOING.
- Write more of payload assembly procedure (Fabian)
- Write more of main assembly procedure (Mark)
- Longer term to-do list
- Do BS Simulink model (Mark, Kokeyama)
- Do 2D drawings of SR2 parts (Hirata)
- Redesign SRx IM wire clamps for larger prisms, remachined RMs.
- Order SRM/PRM parts (Hirata/Shoda)
- Order SRx fasteners (Hirata)
- Type-Bp modification 3D drawings
- Detailed design is ongoing at ATC. (To be updated at 8/5.)
- Installation jig for the protection of the Traverser is under discussion.
- F7 LVDT
- The test bench for the small F7 LVDT will be made at first.
- Sato-san made the drawings for the small test bench. They will be ordered soon.
- Okutomi-kun is starting the test of Virgo F7 LVDT.
- Installation related issues
- Waiting for the update.
- Simulation and others
YAW coupling ratio is estimated from the iKAGRA PR3 data k-log
- Simulation without OSEM sensors and with the length sensing optical sensor.
- Check the existing items and their status
- Although I have not finish ordering procedure, a simplified TAMA TM/RM type 3D model has almost been completed.
- Talking with Prof. Moriwiki (Toyama Univ.) about interference between IMMT1 suspension and intensity stabilization system.
- Found a problem that we need to remove IFI temporarily in order to install IMMT2 from the side window.
Watchdog system (Tanaka)
Materials to use in meeting.PDF
- Organize a meeting with AEL about the interface with the digital system (including the release sensor) (Aso)
- No issue
- Inertial sensor selection for Type-B update (Fabian)
- Wide OSEM
- Mass production
- Deadline for the mass production will be the middle of November.
- Will send a production schedule today (Akutsu)
- Mass production
- Document Tree
- Put links to VIS related documents to the above document tree.
Site work plan for the next week
- Thu: BS retrieval at Kashiwa
- Travel (Week of 8/1):
- Mark: 4 (Kashiwa)
- Hirata: 4 (Kashiwa)
- Travel (Week of 8/8):
Shoda:8-10 (check existing parts)
2016/8/5(Fri) 13:30 -
No F2F in August